a circle with radius 6
Answer by checkley71(8403)

IF YOU WANT THE CIRCUMFERANCE THEN C=2*PI*R OR C=2*3.14*6 OR 37.68
IF YOU WANT THE AREA THEN A=PI*R~2 OR A=3.14*6~2 OR A=3.14*36 OR 113.04